OpenStudy (anonymous):

solve for x. lnx + ln(x-1) = 1

Parth (parthkohli):

\[\ln\left| x(x - 1)\right| = 1\]\[\log_{\large e}x(x - 1) = 1 \\ \\ e = x(x - 1) \]

OpenStudy (jiteshmeghwal9):